OpenStudy (anonymous):

The product of the page numbers on two facing pages of a book is 182. Find the page numbers.

OpenStudy (anonymous):

n(n-1)=182 n=14 so th page is n, n-1=14,13

OpenStudy (blacksteel):

If the two numbers are facing, they will be 1 apart. Thus, we can express them as x and x+1. Then we know that x(x+1) = 182, or x^2 + x - 182 = 0. We can use the quadratic formula to find the roots of this equation, x = -14 and x = 13. Obviously x = -14 doesn't make any sense, so the solution is x = 13, and your pages are 13 and 14.